OpenStudy (adunb8):

linear equation & linear algebra question: i need help with 2nd order differential equations

OpenStudy (adunb8):

\[[d^2x/dt^2 = 2dx/dt + 2]\]

OpenStudy (anonymous):

Wave function?

OpenStudy (lgbasallote):

find x?

OpenStudy (unklerhaukus):

\[\ddot x=2\dot x+2\]\[\ddot x-2\dot x=2\] \[m^2-2m=0\]\[m(m-2)=0\]\[m_{1,2}=0,2\] \[x_c=c_1e^{0x}+c_2e^{2x}\]\[\qquad=c_1+c_2e^{2x}\]

OpenStudy (anonymous):

Ah yes, the characteristic equation. Good call.

OpenStudy (adunb8):

wave function?

OpenStudy (anonymous):

Never mind, thought it looked like x'' = x for a second.

OpenStudy (anonymous):

Ugh, sorry, x''+x=0 is what I was imagining. Sorry it's been a while since I've tried these.

OpenStudy (anonymous):

But, anyway, this was billed as a linear equations/linear algebra question. ?

OpenStudy (unklerhaukus):

\[\ddot x_p-2\dot x_p=2\] \[x_p=c_3x,\qquad \dot x_p=c_3, \qquad\ddot x_p=0,\] \[0-2c_3=2\]\[c_3=-1\] \[x_p=-x,\]

OpenStudy (unklerhaukus):

\[x(t)=x_c+x_p\]

OpenStudy (unklerhaukus):

could you follow what i did @adunb8?

OpenStudy (adunb8):

um... im lost

OpenStudy (unklerhaukus):

which bit,

OpenStudy (adunb8):

what the teacher wants me to do is to turn into 1st order differential equation

OpenStudy (unklerhaukus):

oh, ok, so \(\ddot x=2\dot x+2\) is independent of \(t\) , let \[\dot x=p\] \[\ddot x=p\frac{\text d p}{\text dx}\]

OpenStudy (adunb8):

hm...

OpenStudy (adunb8):

okay thanks i will try out and see if it works out